OpenStudy (yrivers36):

Please help with explanations. I know that I have to square both sides. sqrt of x =sqrt of x-5) +1

OpenStudy (anonymous):

sqrt(x) =sqrt(x-5) +1 sqrt(x) - 1 = sqrt(x-5) square both sides: x - 2sqrtx + 1 = x - 5 -2sqrtx = -6 sqrtx = 3 x = 9 now check this solution if it is a solution..

OpenStudy (anonymous):

i'll explain if you need 'em..

OpenStudy (yrivers36):

im confused with x-2sqrt +1=x-5

OpenStudy (anonymous):

ok...

OpenStudy (anonymous):